matt mixed 6 L of brand A fruit drink and 2 L of brand B fruit drink. which contains 25% fruit juice. FInd the pervent of the fruit juice in brand A if the mixture contained 10% fruit juice
1
Expert's answer
2012-09-27T10:40:19-0400
Let brand A contains X L of juice. Then

X + 2*0.25 = (6 + 2)*0.1

and so, X = 0.3. Therefore, the part of the fruit juice in brand A is

P = 0.3/6 = 0.2

and brand A fruit drink contains 20% of juice.